HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

IR - MILLIMETER-WAVE DOUBLE RESONANCE FOR PRECISION MOLECULAR SPECTROSCOPY

\underline{L. JUPPET} , O. PIRALI, Institut des Sciences Moléculaires d’Orsay, CNRS, Université Paris Saclay, 91405 Orsay; J. F. LAMPIN, Institut d’Electronique

Microélectronique et Nanotechnologie, CNRS, Université de Lille, 59652 Villeneuve d’Ascq cedex; B. DARQUIE, M. MANCEAU, Y. LIU, O. LOPEZ, Laboratoire de Physique des Lasers, CNRS, Université Sorbonne Paris Nord, 93430 Villetaneuse

Metrology-grade spectroscopic measurements of molecules are of high importance for numerous applications such as testing fundamental physics^{1}, molecular physics^{2}, and astrophysics^{3}. In this context, we tested the well-established double resonance technique^{4}^{,}^{5}^{,}^{6}^{,}^{7} using both infrared and millimeter wave transitions to provide high resolution and frequency metrology dataset. The hyperfine components of rotational transitions involving both the ground and the v_{2} = 1 vibrational state of NH_{3} as well as the v_{2} = 1 vibrational state of D_{2}O have been investigated trough IR-THz Doppler-free and cross-over-free double resonance spectroscopy. Spectra were obtained using both a free running external cavity quantum cascade laser (QCL installed at SOLEIL synchrotron) as well as a sub-Hz QCL^{8} (installed at LPL Villetaneuse) frequency-stabilised on the REFIMEV signal^{9}. The lasers have been used to optically pump the molecules inducing a velocity class selection. We will present the major interests of this approach for precision spectroscopy and the current limitations we meet.

1M. R. Fiecher and al, The Journal of Physical Chemistry Letters 13 42 (2022)

2S. P. Belov and al, The Journal of Chemical Physics 145 (2016)

3C. Puzzarini and al, PCCP 12 (2020)

4J. Lemaire, Thèse de l’Université de Lille (1974)

5T. Shimizu and al, The Journal of chemical Physics 53 (1970)

6S. Ghoshal and al, The Journal of chemical Physics 83 (1985)

7A. Foltynowicz and al, PRL 126 (2021)

8R. Santagata and al, Optica 6 (2019)

9https://www.refimeve.fr p-number: p245  Submitted on Mon, 26 Jun 2023 15:07:27 +0200